# Analysis of Fourier Neural Operators via Effective Field Theory

### v2

@@ 4.1 Analytic Activations
$\begin{split}&\mathcal{K}^{(l+1)}(f,f^{\prime})=\delta(f-f^{\prime})C_{R^{(l+1)}}(f)\sum_{k=1}^{\infty}\frac{\sigma_{k}^{2}}{(n^{(l)})^{k-1}}\\ &\sum_{n_{k,1}+\dots+n_{k,l}=k}\frac{(n_{k,1})!\dots(n_{k,l})!}{k!}(\mathcal{H}^{(l)})^{*n_{k,1}}\cdots(\mathcal{H}^{(l)})^{*n_{k,l}}(f)\\ &\chi_{\parallel}(f,f^{\prime})=\delta(f-f^{\prime})C_{R^{(l+1)}}(f)\sum_{k=1}^{\infty}\frac{\sigma_{k}^{2}}{(n^{(l)})^{k-1}}\\ &\sum_{n_{k,1}+\dots+n_{k,l}=k}\frac{(n_{k,1})!\dots(n_{k,l})!}{(k-1)!}\frac{(\mathcal{H}^{(l)})^{*n_{k,1}}\cdots(\mathcal{H}^{(l)})^{*n_{k,l}}(f)}{\mathcal{K}^{(l)}\{\mathbf{u_{0}}\}(f)},\\ &\tilde{\chi}_{\perp}(f,f^{\prime})=\sum_{k=1}^{\infty}\frac{\sigma_{k}^{2}}{(n^{(l)})^{k-1}}\\ &\sum_{n_{k,1}+\dots+n_{k,l}=k-1}\frac{(n_{k,1})!\dots(n_{k,l})!}{(k-1)!}(\mathcal{H}^{(l)})^{*n_{k,1}}\cdots(\mathcal{H}^{(l)})^{*n_{k,l}}(f).\end{split}$
4.2 Scale Invariant Activations
The scale invariant class of activations is defined as follows:
@@ 4.2 Scale Invariant Activations
Theorem 2 With the scale invariant activation specified in Eq. 15 and a Fourier Neural Operator defined by Eq. 4 under the initialization ensemble Eq. 5, the kernel and the susceptibilities are given by the following recursion relations:
@@ 4.3 Architecture with Residual Connection
In practice, an FNO keeps only a subset of Fourier modes: after the FFT, all modes above a cutoff are truncated. With no activation (purely linear layers), both pre-activations and outputs stay strictly within this band. Once a nonlinearity is added, however, low frequencies couple to high frequency kernels—as shown in Sections 4.1–4.2 and confirmed empirically in Appendix H. For inputs of modest extent and large channel width, Theorem 1 implies that this coupling is weak. To retain the efficiency of truncation while still accumulating nonlinear interactions, we adopt a ResNet style update that adds each layer’s pre-activation to its activation output. This helps preserve high frequency information with depth and improves fine scale processing. The rest of this section develops the corresponding theory for residual FNOs. We begin by defining a weighted residual connection for the kernel integration layer, modifying Eq. 4 accordingly.
@@
$\begin{split}&\mathcal{K}^{(l+1)}\{\mathbf{u}\}(f,f^{\prime})=\delta(f-f^{\prime})\Big{(}(\sigma_{1}^{2}C_{R^{(l+1)}}(f)+\gamma)\mathcal{K}^{(l)}\{\mathbf{u}\}(f)\\ &+C_{R^{(l+1)}}(f)\sum_{k=2}^{\infty}\frac{\sigma_{k}^{2}}{(n^{(l)})^{k-1}}\sum_{n_{k,1}+\dots+n_{k,l}=k}\frac{(n_{k,1})!\dots(n_{k,l})!}{k!}(\mathcal{H}^{(l)})^{*n_{k,1}}\cdots(\mathcal{H}^{(l)})^{*n_{k,l}}(f)\Big{)}.\\ &\chi_{\parallel}(f,f^{\prime})=\delta(f-f^{\prime})\Big{(}\gamma+C_{R^{(l+1)}}(f)\sum_{k=1}^{\infty}\frac{\sigma_{k}^{2}}{(n^{(l)})^{k-1}}\\ &\sum_{n_{k,1}+\dots+n_{k,l}=k}\frac{k(n_{k,1})!\dots(n_{k,l})!}{(k-1)!}\frac{(\mathcal{H}^{(l)})^{*n_{k,1}}\cdots(\mathcal{H}^{(l)})^{*n_{k,l}}(f)}{\mathcal{K}^{(l)}\{\mathbf{u_{0}}\}(f)}\Big{)}.\\ &\tilde{\chi}_{\perp}(f,f^{\prime})=\Big{(}\sum_{k=1}^{\infty}\frac{\sigma_{k}^{2}}{(n^{(l)})^{k-1}}\\ &\sum_{n_{k,1}+\dots+n_{k,l}=k-1}\frac{(n_{k,1})!\dots(n_{k,l})!}{(k-1)!}(\mathcal{H}^{(l)})^{*n_{k,1}}\cdots(\mathcal{H}^{(l)})^{*n_{k,l}}(f)\Big{)}+(\gamma+2\tilde{\gamma}\sigma_{1})\delta(f)\delta(f^{\prime}).\end{split}$
@@ 5.1 Kernel Evolution Prediction
In this subsection we start by examining how the kernel evolves under three different activations—quadratic, tanh, and ReLU—across a range of network depths and widths. We then study how the kernel changes when a tanh activation is used inside a ResNet-style FNO. Secondly, using the kernel recursion formula derived in Section 4, Appendix G, we feed the empirical kernel from layer $l$ into the recurrence to predict the kernel at layer $l+1$ and compare that prediction with the measured kernel. To validate the frequency coupling and theoretical prediction over the truncation frequency ($k_{\text{max}}$), we consider the reduced kernel $\tilde{\mathcal{K}}$ which is defined as follows:
@@
In the post-truncation band $f\in[17,63]$ ($k_{max}=16$), the measured kernels closely follow our theoretical curves across activation families (tanh, cubic, ReLU) and with/without a residual connection. With channel width fixed at 32 and depths 0,2 for tanh/cubic and 0,1,2,3 for ReLU/residual, the theory lies almost entirely within the $\pm 1$ s.d. band over 100 runs. Ablation tests that remove individual activation terms break this agreement, indicating that all relevant terms must be retained in the theory. The same behavior holds for non-constant spectra, e.g., $C_{R}(f)=\frac{2.5}{|f|}$. See Appendix H, Figs. 2 and 3 for full details.
@@
Figure 1: Empirical vs. theoretical susceptibilities (log scale; red error bars = $\pm 1$ s.d., $N=100$) for width $n=32$ and truncation $k_{\max}=32$. Row 1: parallel susceptibility $\chi_{\parallel}$ for (col 1) quadratic with step $C_{R}$, (col 2) cubic with step $C_{R}$, (col 3) cubic with $C_{R}(f)\propto 1/|f|$ (truncated). Row 2: corresponding perpendicular susceptibility $\chi_{\perp}$. Row 3: $\chi_{\parallel}$ for cubic + residual with $\gamma=0.25$ (col 1) and $\gamma=0.75$ (col 2), and ReLU (col 3), all with step $C_{R}$. Row 4: corresponding $\chi_{\perp}$. Parallel runs use $\varepsilon_{\parallel}=0.01$; perpendicular runs use $\varepsilon_{\perp}=10^{-5}$. Markers = measurements; solid curves = theory.

### v3

@@ 4.1 Analytic Activations
$\begin{split}&\mathcal{K}^{(l+1)}(f,f^{\prime})=\delta(f-f^{\prime})C_{R^{(l+1)}}(f)\sum_{k=1}^{\infty}\frac{\sigma_{k}^{2}}{(n^{(l)})^{k-1}}\\ &\sum_{n_{k,1}+\dots+n_{k,l}=k}\frac{(n_{k,1})!\dots(n_{k,l})!}{k!}(\mathcal{H}^{(l)})^{*n_{k,1}}\cdots(\mathcal{H}^{(l)})^{*n_{k,l}}(f),\\ &\chi_{\parallel}(f,f^{\prime})=\delta(f-f^{\prime})C_{R^{(l+1)}}(f)\sum_{k=1}^{\infty}\frac{\sigma_{k}^{2}}{(n^{(l)})^{k-1}}\\ &\sum_{n_{k,1}+\dots+n_{k,l}=k}\frac{(n_{k,1})!\dots(n_{k,l})!}{(k-1)!}\frac{(\mathcal{H}^{(l)})^{*n_{k,1}}\cdots(\mathcal{H}^{(l)})^{*n_{k,l}}(f)}{\mathcal{K}^{(l)}\{\mathbf{u_{0}}\}(f)},\\ &\tilde{\chi}_{\perp}(f,f^{\prime})=\delta(f-f^{\prime})\sum_{k=1}^{\infty}\frac{\sigma_{k}^{2}}{(n^{(l)})^{k-1}}\\ &\sum_{n_{k,1}+\dots+n_{k,l}=k-1}\frac{(n_{k,1})!\dots(n_{k,l})!}{(k-1)!}(\mathcal{H}^{(l)})^{*n_{k,1}}\cdots(\mathcal{H}^{(l)})^{*n_{k,l}}(f).\end{split}$
4.2 scale-invariant Activations
The scale-invariant class of activations is defined as follows:
@@
Theorem 2 With the scale-invariant activation specified in Eq. 15 and a Fourier Neural Operator defined by Eq. 4 under the initialization ensemble Eq. 5, the kernel and the susceptibilities are given by the following recursion relations:
@@ 4.3 Architecture with Residual Connection
In practice, an FNO keeps only a subset of Fourier modes: after the FFT, all modes above a cutoff are truncated. This truncation can exacerbate spectral bias and hinder the learning of fine-scale features; multi-resolution wavelet-based constructions alleviate this by enriching the representation across scales (Han et al. (2026)), and complementary strategies explicitly promote high-frequency learning in neural operators (Khodakarami et al. (2026)). With no activation (purely linear layers), both pre-activations and outputs stay strictly within this band. Once a nonlinearity is added, however, low frequencies couple to high frequency kernels—as shown in Sections 4.1–4.2 and confirmed empirically in Section 5.1. For inputs of modest extent and large channel width, Theorem 1 implies that this coupling is weak. To retain the efficiency of truncation while still accumulating nonlinear interactions, we adopt a ResNet style update that adds each layer’s pre-activation to its activation output. This helps preserve high frequency information with depth and improves fine scale processing. The rest of this section develops the corresponding theory for residual FNOs. We begin by defining a weighted residual connection for the kernel integration layer, modifying Eq. 4 accordingly.
@@
$\begin{split}&\mathcal{K}^{(l+1)}\{\mathbf{u}\}(f,f^{\prime})=\delta(f-f^{\prime})\Big((\sigma_{1}^{2}C_{R^{(l+1)}}(f)+\gamma)\mathcal{K}^{(l)}\{\mathbf{u}\}(f)\\ &+C_{R^{(l+1)}}(f)\sum_{k=2}^{\infty}\frac{\sigma_{k}^{2}}{(n^{(l)})^{k-1}}\sum_{n_{k,1}+\dots+n_{k,l}=k}\frac{(n_{k,1})!\dots(n_{k,l})!}{k!}(\mathcal{H}^{(l)})^{*n_{k,1}}\cdots(\mathcal{H}^{(l)})^{*n_{k,l}}(f)\Big),\\ &\chi_{\parallel}(f,f^{\prime})=\delta(f-f^{\prime})\Big(\gamma+C_{R^{(l+1)}}(f)\sum_{k=1}^{\infty}\frac{\sigma_{k}^{2}}{(n^{(l)})^{k-1}}\\ &\sum_{n_{k,1}+\dots+n_{k,l}=k}\frac{k(n_{k,1})!\dots(n_{k,l})!}{(k-1)!}\frac{(\mathcal{H}^{(l)})^{*n_{k,1}}\cdots(\mathcal{H}^{(l)})^{*n_{k,l}}(f)}{\mathcal{K}^{(l)}\{\mathbf{u_{0}}\}(f)}\Big),\\ &\tilde{\chi}_{\perp}(f,f^{\prime})=\delta(f-f^{\prime})\Big(\sum_{k=1}^{\infty}\frac{\sigma_{k}^{2}}{(n^{(l)})^{k-1}}\\ &\sum_{n_{k,1}+\dots+n_{k,l}=k-1}\frac{(n_{k,1})!\dots(n_{k,l})!}{(k-1)!}(\mathcal{H}^{(l)})^{*n_{k,1}}\cdots(\mathcal{H}^{(l)})^{*n_{k,l}}(f)\Big)+(\gamma+2\tilde{\gamma}\sigma_{1})\delta(f)\delta(f^{\prime}).\end{split}$
@@ 5 Experimental Validations
5.1 Experimental Setup and Data Generation
Input functions.
Our theory conditions on fixed input functions $\mathbf{u}$ and $\mathbf{v}$ and takes expectations only over random parameter initialization. To instantiate $\mathbf{u}$ and $\mathbf{v}$ in experiments, we use a synthetic mean-zero Gaussian random field (GRF) as a convenient testbed: we draw $\mathbf{u}$ (and $\mathbf{v}$ when needed) once from a stationary covariance $C(x,y)=\exp(-\|x-y\|^{2}/\ell^{2})$ (unless stated otherwise), and keep the sampled inputs fixed while averaging over $N=100$ independent random initializations. This separation mirrors the theoretical ensemble and isolates the frequency-transfer effect induced by the network nonlinearity rather than by input variability.
Domain and discretization.
We work on a periodic domain of length $L$ discretized on an equispaced grid with $N_{x}$ points. GRF samples are generated in Fourier space with the corresponding power spectrum and transformed back via FFT, enforcing conjugate symmetry to obtain real-valued signals. Unless otherwise stated, inputs are normalized to have unit $L^{2}$-norm before being fed to the network.
Reproducibility.
@@ 5.2 Kernel Evolution Prediction
Overview.
We study how the layer kernel evolves under three activation families—quadratic, $\tanh$, and ReLU—across varying depths and widths, and then examine a ResNet-style FNO with a $\tanh$ nonlinearity. Beyond qualitative inspection of frequency transfer, we also test the theory quantitatively: given an empirical kernel at layer $l$, we use the recursion derived in Section 4 and Appendix G to predict the kernel at layer l+1l\!+\!1, and compare this prediction against measurements under independent random initializations.
Reduced kernel and post-truncation evaluation.
To isolate the mechanism of frequency coupling beyond the truncation wavenumber $k_{\max}$, we work with the reduced kernel
@@
i.e., the Gaussian expectation term that appears in the kernel update prior to multiplication by the spectral profile $C_{R}$. We place particular emphasis on the post-truncation band (e.g., $f\in[17,63]$ for $k_{\max}=16$), where purely linear truncated Fourier layers would yield identically zero energy.
Qualitative frequency transfer: linear vs. nonlinear layers.
Figs. 2 to 5 compare networks with and without nonlinear activation under a step profile $C_{R}(f)=\mathbf{1}_{\{|f|\leq k_{\max}\}}$. We consider two width–depth pairs, $n\in\{4,64\}$ and $L\in\{2,4\}$, with inputs sampled from a mean-zero Gaussian random field with covariance $C(x,y)\propto\exp(-\|x-y\|^{2}/l^{2})$ and truncation $k_{\max}=16$. For each configuration, we initialize $N=100$ independent models and plot (in log scale) the kernel magnitude as a function of frequency: dots denote individual runs, while the thick curve shows the ensemble mean. For the quadratic activation $\sigma(x)=x^{2}+x$ (Fig. 2), the kernel drops to numerical zero immediately after $k_{\max}$ when the activation is removed, reflecting strict band limitation in the linear truncated operator. With the nonlinearity present, nonzero energy persists up to approximately $2k_{\max}$, consistent with the leading double-convolution term in the analytic recursion. For $\tanh$ (Fig. 3), the Taylor-series contributions populate substantially higher wavenumbers, so frequency transfer remains visible well beyond $2k_{\max}$. For ReLU (Fig. 4), the decay beyond $k_{\max}$ is markedly slower even without residual connections, matching the scale-invariant theory: the update is governed by the spatial correlation $\rho^{(l)}(x,x^{\prime})$, whose sharp structure in position space induces broad spectral support.
Residual connections and depth-wise retention.
Figure 5 further shows the effect of residual updates with $\tanh$. To prevent kernel blow-up, the weight variance is scaled by $1-\gamma$. Small $\gamma$ yields rapid post-truncation decay with depth, whereas large $\gamma$ preserves substantial high-frequency energy, in line with the residual criticality analysis in Section 4.3.
Width scaling and finite-width fluctuations.
Two consistent trends emerge across Figs. 2 to 5. First, increasing width concentrates the empirical kernels around the ensemble mean, consistent with the theoretical $O(1/n)$ suppression of the four-point vertex (and hence of kernel fluctuations). Second, larger widths also reduce the absolute magnitude of post-truncation energy, as predicted by the width-suppressed higher-order terms in the analytic recursion.
Quantitative agreement with the kernel recursion.
@@
We quantitatively test the kernel recursion by a one-step prediction protocol: given the empirical kernel at layer $l$, we apply the theoretical update (Section 4 and Appendix G) to predict the kernel at layer l+1l\!+\!1 and compare it against the measured kernel under independent random initializations. In the post-truncation band $f\in[17,63]$ for $k_{\max}=16$, the measured reduced kernels closely follow the theoretical curves across activation families (tanh, cubic, ReLU) and with/without residual connections. Fixing the width to $n=32$, we report depths $l\in\{0,2\}$ for tanh/cubic and $l\in\{0,1,2,3\}$ for ReLU/residual; in almost all results the theoretical curve lies within the empirical mean $\pm 1$ std band over $N=100$ runs (Figs. 6 to 9). Ablations that remove individual activation terms from the theoretical update systematically break this agreement, indicating that all relevant contributions must be retained. The same level of agreement persists for non-constant truncated spectra, e.g. $C_{R}(f)=2.5/|f|$, both with and without residual connections (Fig. 9).
Figure 2: Reduced kernel (log scale) quadratic, no residuals, widths $n\in\{4,64\}$ at depths $l=2$ (top) and $l=4$ (bottom).
Figure 3: Reduced kernel (log scale) tanh, no residuals, widths $n\in\{4,64\}$ at depths $l=2$ (top) and $l=4$ (bottom).
Figure 4: Reduced kernel (log scale) ReLU, no residuals, widths $n\in\{4,64\}$ at depths $l=2$ (top) and $l=4$ (bottom).
Figure 5: Reduced kernel (log scale) tanh, with residuals, width $n=32$ with $\gamma=0.1,0.9$ at $l=2$ (top) and $l=6$ (bottom).
Figure 6: Empirical measurement vs. theoretical prediction of reduced kernels (log scale) (width $n=32$, truncation $k_{\max}=16$). $\tanh$ (top), cubic (bottom) with step profile $C_{R}(f)=\mathbf{1}_{\{|f|\leq k_{\max}\}}$ $l=0$ (left), $l=2$ (right). Pink shading: mean $\pm 1$ std over $N=100$ initializations; solid curve: theory; dots: measurements.
Figure 7: Empirical measurement vs. theoretical prediction of reduced kernels (log scale) (width $n=32$, truncation $k_{\max}=16$). ReLU with step profile $C_{R}(f)=\mathbf{1}_{\{|f|\leq k_{\max}\}}$ from $l=0$ (top left) to $l=3$ (bottom right). Pink shading: mean $\pm 1$ std over $N=100$ initializations; solid curve: theory; dots: measurements.
Figure 8: Empirical measurement vs. theoretical prediction of reduced kernels (log scale) (width $n=32$, truncation $k_{\max}=16$). Cubic activation with residual connection ($\gamma=0.05$) and step profile $C_{R}(f)=\mathbf{1}_{\{|f|\leq k_{\max}\}}$ from $l=0$ (top left) to $l=3$ (bottom right). Pink shading: mean $\pm 1$ std over $N=100$ initializations; solid curve: theory; dots: measurements.
Figure 9: Empirical measurement vs. theoretical prediction of reduced kernels (log scale) (width $n=32$, truncation $k_{\max}=16$). Cubic activation without (top) and with (bottom) residual connection ($\gamma=0.05$) and step profile $C_{R}(f)\propto 1/|f|$ (truncated at $k_{\max}$) $l=0$ (left), $l=2$ (right). Pink shading: mean $\pm 1$ std over $N=100$ initializations; solid curve: theory; dots: measurements.
@@
Figure 10: Empirical vs. theoretical parallel susceptibility $\chi_{\parallel}$ (log scale; red error bars = $\pm 1$ std, $N=100$) for width $n=32$ and truncation $k_{\max}=32$. Row 1: quadratic and cubic activations with step $C_{R}$. Row 2: cubic + residual with $\gamma=0.25$ and $\gamma=0.75$. Row 3: ReLU with step $C_{R}$ and cubic with $C_{R}(f)\propto 1/|f|$ (truncated). Used $\varepsilon_{\parallel}=0.01$; Markers = measurements; solid curves = theory.
Figure 11: Empirical vs. theoretical perpendicular susceptibility $\chi_{\perp}$ (log scale; red error bars = $\pm 1$ std, $N=100$) for width $n=32$ and truncation $k_{\max}=32$. Row 1: quadratic and cubic activations with step $C_{R}$. Row 2: cubic + residual with $\gamma=0.25$ and $\gamma=0.75$. Row 3: ReLU with step $C_{R}$ and cubic with $C_{R}(f)\propto 1/|f|$ (truncated). Used $\varepsilon_{\perp}=10^{-5}$; Markers = measurements; solid curves = theory.